OpenStudy (anonymous):

determine the values of b and x so that the following function is continuous on the entire line f(x){x+1 1

OpenStudy (anonymous):

sorry typo determine the values of b and c

OpenStudy (anonymous):

is it \[f(x) = \left\{\begin{array}{rcc} x+1 & \text{if} & 1<x<3 \\ x^2+bx+c& \text{if} & |x-2|\geq 1 \end{array} \right. \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so it is easier than you think make sure the two functions agree at \(x=1\) and at \(x=3\)

OpenStudy (anonymous):

what do you mean? make them equal or that they are continuous

OpenStudy (anonymous):

at \(x=1\) you get \(1+1=2\) for the first expression and \(1+a+b\) for the second, so you know \[1+a+b=2\] i.e. \[a+b=1\]

OpenStudy (anonymous):

at \(x=3\) you get \(3+1=4\) for the top one and \[9+3a+b\] for the bottom one, so you know \[4=9+3a+b\] or \[3a+b=-5\]

OpenStudy (anonymous):

how do you know 1+a+b=2

OpenStudy (anonymous):

sorry i used \(a\) and \(b\) instead of \(b\) and \(c\) but no matter, solve \[b+c=1\] \[3b+c=-5\] and you will have your answer

OpenStudy (anonymous):

because if you use the top formula \(x+1\) if you replace \(x\) by \(1\) you get \(2\)

OpenStudy (anonymous):

all you are really trying to do is make the two expressions agree at 1 and 3 where it changes definition that is the whole idea behind the problem

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

the two expression are \(x+1\) and \(x^2+bx+c\) if you replace \(x\) by \(1\) in both and set them equal you get \(1+1=1+b+c\)

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yw